Translate to equation(s) only (do not solve): John has a mean (average) of 84 for 2 tests, and the range is 16. What are the two scores?
Answer by Theo(13342)

let a be the larger nunmber.
let b be the smaller number.

the average is 84.

you get (a + b) / 2 = 84

multiply both sides of this equation by 2 and you get a + b = 168.

the range is 16.

you get a - b = 16.

solve these 2 equations simultaneously and you should get your answer.